OpenStudy (anonymous):

({x-8}/{x-7}) <=-3 FInd the interval i added 3 and tried to solve it but the systems marks me wrong.. i came up with (7,7.25]U[7.25,infinity)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

\[\frac{x-8}{x-7}\le-3~~~~~~~~~~~\text{(right away you know that }x\not=7)\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\text{(you also know that if }x<7,\text{ then the left side is positive})\\ x-8\le-3(x-7)\\ x-8\le-3x+21\\ 4x\le29\\ x\le\frac{29}{4}\] So it looks like you should be getting \((7,7.25]\).

OpenStudy (anonymous):

damn, i was close, thanks for your response

OpenStudy (anonymous):

You're welcome!